OpenStudy (anonymous):

Find the GCF of 27 and 45. 45 27 9 3

OpenStudy (anonymous):

9. 27 and 45 can both be divided by 9 and is the biggest number that can.

OpenStudy (anonymous):

45=27*1+18 27=18*1+9 18=9*2+0 G.C.F=9